# Pinning control of chimera states in systems with higher-order interactions

### v1

@@ III.1 Comparison between higher-order and pairwise interactions
Let us conclude by pointing out that in previous works chimera states were obtained for specific values of the initial conditions, while random or uniform initial conditions did not yield the same result. In our numerical study, the initial conditions do not matter as long as the system starts on the synchronous solution, i.e., on the limit cycle of the Stuart-Landau oscillator (1). Hence, we will choose the initial conditions to be uniform and without noise, i.e., $(x_{j}^{0}=1,y_{j}^{0}=0)$ for every oscillator.
@@ III.2 Scaling of the pinned subset with respect to the hyperring size
It is already remarkable that the higher-order framework favors the onset of chimera states with respect the pairwise one, but here we also show that the percentage of nodes required to obtain a chimera state, by exploiting the interactions of the hyperring, is small. Hence, we set up a pinning protocol in which we try to maximize the number of nodes affected by one single controller. Due to the hyperring structure, a way can be to pin every other junction node, so that each controlled node can, in principle, affect two hyperedges, as schematically shown in Fig. 7. To observe how the number of pinned nodes scales with the size of the hyperring, we keep constant the number of hyperedges, so that the total number of nodes increases, but the backbone of the structure remains unchanged. Given that each pinned node is part of $2$ hyperedges, in principle, we are able to control all the nodes in these hyperedges. E.g., in a $3$-hyperring, with each pinned nodes we would control $7$ nodes, in a $4$-hyperring $9$ nodes, in a $5$-hyperring $11$ nodes and in a $6$-hyperring $13$ nodes. For brevity, we present here the results for the additive pinning. The results for the parametric pinning are similar and can be found in Appendix B.
@@
Figure 8: Scaling of the pinned subset with respect to the higher-order structure size with the additive pinning approach. On the upper panels we depict the snapshots of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units, while the lower panels show the hyperedge-based local order parameters. Panels a) and b) report the simulations performed on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes, panels c) and d) on a $4$-hyperring (i.e., $5$-body interactions) of $272$ nodes, panels e) and f) on a $5$-hyperring (i.e., $6$-body interactions) of $340$ nodes and panels g) and h) on a $6$-hyperring (i.e., $7$-body interactions) of $408$ nodes. The number of nodes is chosen so that each hyperring is made of $68$ hyperedges. Pinning control is applied to $N_{p}=18$ nodes as in Fig. 7, i.e., every two junction nodes, to all the structures and for $t_{p}=100$ time units. This means that the pinned nodes are $\simeq 8.8\%$ of the total nodes in the $3$-hyperring, $\simeq 6.6\%$ of the total nodes in the $4$-hyperring, $\simeq 5.3\%$ of the total nodes in the $5$-hyperring and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring. The model parameters are $\alpha=1$ and $\omega=1$, the coupling strength is $\varepsilon=0.01$, except for the $6$-hyperring where $\varepsilon=0.1$, and the parameters $\lambda_{i_{p}}$ are drawn from a uniform distribution in the interval $[0,2]$ and are the same for all the simulations.
@@
In Fig. 8, we show the results obtained with such control scheme on $d$-hyperrings with $d=3,4,5,6$, i.e., $4$-, $5$-, $6$- and $7$-body interactions, where we have fixed the number of hyperedges. Indeed, we can observe that we obtain a chimera state by inducing a large region of incoherence (more than half of the nodes) with a control that involves only a small fraction of the nodes. Moreover, the pinned nodes $N_{p}$ are kept constant for every structure, meaning that $N_{p}$ does not scale with the number of nodes, but rather with the number of hyperedges, which allows to control large structures with only a handful of nodes. The pinned nodes are $\simeq 8.8\%$ of the total nodes for the $3$-hyperring (panels a) and b)), $\simeq 6.6\%$ of the total nodes for the $4$-hyperring (panels c) and d)), $\simeq 5.3\%$ of the total nodes in the $5$-hyperring (panels e) and f)) and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring (panels g) and h)). In the latter case, our pinning scheme does not work as well as for the other structures and the coupling strength need to be significantly increased in order to observe a chimera state. However, we can observe from Fig. 8h) that the chimera is not anymore stable and the front of incoherence enlarges. Let us stress that stable chimera states can be easily obtained also in $6$-hyperrings by reducing the distance between the controlled nodes, as in the previous section, instead of the pinning scheme of Fig. 7.
@@
Let us note that the parameters $\lambda_{i_{p}}$ need to be all positive or all negative in order for this pinning scheme to yield persistent chimera states. When such random inputs are drawn in a symmetric interval with respect to the 00, e.g., $[-2,2]$ as in Fig. 3, a chimera states forms but vanishes at about $400$ time units. On the other hand, when the control inputs have the same sign, e.g., $[0,2]$ as in Fig. 8, the chimera is persistent until about $4000$ time units, i.e., $10$ times longer.
@@
In Appendix B, we show analogous results for the parametric pinning, except for the case $6$-hyperring, where chimera states do not emerge by pinning every other junction node, even with stronger couplings. The fact that our results are robust with respect to different control approaches is a good indication of the pervasiveness of the phenomenon.
@@ IV Heuristic interpretation through phase reduction theory
In this section we will give a heuristic interpretation of the results based on the phase reduction approach Nakao (2016); Kuramoto and Nakao (2019), which consists of reducing a given oscillatory system to a phase, i.e., Kuramoto-type, model Kuramoto (1975). In a nutshell, starting from a system of $N$ highly dimensional units in a limit cycle regime, e.g.,
@@
under given assumptions (see Nakao (2016) for details), we can reduce to a system of $N$ interacting phase oscillators of the form
$\dot{\vartheta}_{i}=\omega_{i}+\varepsilon\sum_{j=1}^{N}A_{i,j}\vec{Z}(\vartheta_{i})\cdot\vec{G}_{ij}(\vartheta_{j},\vartheta_{i}),$ (16)
where $\omega_{i}$ is the frequency of the $i$-th oscillator and $\vec{Z}$ is the phase sensitivity function. The key in the reduction process is to find an expression for $\vec{Z}$, which has an analytical expression only for some specific cases, among which the Stuart-Landau model Nakao (2016) and weakly nonlinear oscillators León and Nakao (2023), while it needs to be obtained numerically for general oscillators.
The phase reduction theory has been applied also to systems with higher-order interactions Ashwin and Rodrigues (2016); León and Pazó (2019); León et al. (2024), obtaining higher-order versions of the Kuramoto model, which exhibit much richer behaviors than the pairwise one. Indeed, the first evidence of higher-order-induced exotic behaviors, which triggered the interest of the community towards this new framework, has come from higher-order phase models (although not derived through phase reduction) Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020). In what follows, we will apply the phase reduction to our model on a $3$-hyperring and on the corresponding clique-projected network, Eqs. (6) and (7) respectively, to give an intuition of why it is easier to induce chimera behavior via pinning when the topology is higher-order.
For the Stuart-Landau model, the phase sensitivity function is $\vec{Z}(\vartheta)=(-\sin(\vartheta),\cos(\vartheta))^{\top}$ Nakao (2016). Let us consider system (6) in polar coordinates, i.e., $\vec{X}_{i}=(\cos(\vartheta_{i}),\sin(\vartheta_{i}))$, and proceed with the reduction by computing the following
@@
$-\alpha\cos(\vartheta_{i})\sin(\vartheta_{i})+\omega\sin^{2}(\vartheta_{i})+\cos(\vartheta_{i})\sin(\vartheta_{i})+\omega\cos^{2}(\vartheta_{i})+\alpha\cos(\vartheta_{i})\sin(\vartheta_{i})-\cos(\vartheta_{i})\sin(\vartheta_{i})+$
$\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{i,j_{1},j_{2},j_{3}}^{(3)}\Big{(}\cos^{3}(\vartheta_{i})\sin(\vartheta_{i})-\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\sin(\vartheta_{i})+\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\cos(\vartheta_{i})-\cos^{4}(\vartheta_{i})\Big{)},$
@@
$\dot{\vartheta}_{i}=\omega+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{i,j_{1},j_{2},j_{3}}^{(3)}\Phi(\vartheta_{i},\vartheta_{j_{1}},\vartheta_{j_{2}},\vartheta_{j_{3}}),$ (17)
@@
$\Phi(\vartheta_{i},\vartheta_{j_{1}},\vartheta_{j_{2}},\vartheta_{j_{3}})=\cos^{3}(\vartheta_{i})\sin(\vartheta_{i})-\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\sin(\vartheta_{i})+\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\cos(\vartheta_{i})-\cos^{4}(\vartheta_{i}).$ (18)
@@
$\dot{\vartheta}_{i}=\omega+\varepsilon\sum_{j}^{N}A_{i,j}\Psi(\vartheta_{i},\vartheta_{j}),$ (19)
@@
Through the averaging method Nakao (2016), the $4$-body coupling (18) can be approximated as
@@
The coupling given by Eq. (22) steers the system towards synchronization, because $\Psi(\vartheta_{i},\vartheta_{j})$ has form of the well-known Kuramoto-Sakaguchi coupling, i.e., $\sin(\vartheta_{j}-\vartheta_{i}+\alpha)$, which is known to be attractive for $|\alpha|<\frac{\pi}{2}$ Kuramoto (1984). On the other hand, Eq. (21) allows for a much richer dynamics, given that there are many more combinations of the phases and the coefficients such that the coupling term vanishes, as it is the case of the higher-order Kuramoto model León et al. (2024). This fact, gives an intuition not only of the much richer dynamics observed when higher-order interactions are present Tanaka and Aoyagi (2011); Skardal and Arenas (2019, 2020); Millán et al. (2020); León et al. (2024), but also of why higher-order systems favor the presence of chimera states and it is much easier to induce such behavior via pinning, compared to the pairwise case. Moreover, given the form of the higher-order coupling terms, once such state is achieved, it is more difficult for the higher-order system to steer towards synchronization, which provides a qualitative explanation of why the chimera states are also persistent. Let us remark that the first intuition of this behavior was given by Zhang et al. in Zhang et al. (2023), where it was shown, for the higher-order Kuramoto model, that, when the systems leaves the attraction basin of the synchronous state, it is more difficult to synchronize again because higher-order interactions cause a shrinking of such attraction basin, which becomes smaller. In our case, the control pushes the system away from the synchronous solution creating a chimera state and the higher-order interactions favor the persistence of this state.
@@
Figure 9: Additive pinning induces phase chimera states for the reduced phase model (17) on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of the phases $\vartheta_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$. The frequency is $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of Figs. 3 and 4.
@@
Figure 10: Additive pinning for the reduced phase model (19) on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 9. Panel a) depicts the whole time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of the phases $\vartheta_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$. The frequency is $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of Figs. 3 and 4.
@@
As a further corroboration of the results shown in this work and of the correctness of our phase reduction approach, let us conclude this section by applying our pinning protocols to the reduced phase models and show that the outcome confirms our claims. We will show the results of additive pinning (10), leaving those for the parametric approach (12) in Appendix C. In Figs. 9 and 10, we show the result of the additive pinning procedure on a $3$-hyperring and its corresponding clique-projected network, respectively. The setting is analogous to that of Figs. 3 and 4, and so are the results: while the control induces a chimera state in the higher-order phase reduced model, no chimera emerges in the pairwise setting. The same results are obtained also when performing parametric pinning control (see Appendix C).

### v2

@@ III.1 Comparison between higher-order and pairwise interactions
Let us conclude by pointing out that, in previous works, chimera states were obtained for specific values of the initial conditions, while random or uniform initial conditions did not yield the same result. In our numerical study, the initial conditions do not matter as long as the system starts on the synchronous solution, i.e., on the limit cycle of the Stuart-Landau oscillator (1). Hence, we will choose the initial conditions to be uniform and without noise, i.e., $(x_{j}^{0}=1,y_{j}^{0}=0)$ for every oscillator.
@@ III.2 Scaling of the pinned subset with respect to the hyperring size
The previous setting of consecutively pinned nodes exhibit some features that are independent of the size but are caused by the sole structure of the topology. In fact, independently of the total size of the hyperring and of the number of pinned nodes, the nodes which are affected by the perturbation outside of the pinning region are also of the order of $2$ hyperedges. Hence, the chimera state will develop on the pinned region with a boundary of about two hyperedges. We can exploit this property of the topology to obtain chimera states with a very small percentage of pinned nodes. Hence, we set up a pinning protocol in which we try to maximize the number of nodes affected by one single controller. Due to the hyperring structure, a way can be to pin every other junction node, so that each controlled node can, in principle, affect two hyperedges, as schematically shown in Fig. 7. To observe how the number of pinned nodes scales with the size of the hyperring, we keep constant the number of hyperedges, so that the total number of nodes increases, but the backbone of the structure remains unchanged. Given that each pinned node is part of $2$ hyperedges, in principle, we are able to control all the nodes in these hyperedges. E.g., in a $3$-hyperring, with each pinned nodes we would control $7$ nodes, in a $4$-hyperring $9$ nodes, in a $5$-hyperring $11$ nodes and in a $6$-hyperring $13$ nodes. For brevity, we present here the results for the additive pinning. The results for the parametric pinning are similar and can be found in Appendix B.
@@
Figure 8: Scaling of the pinned subset with respect to the higher-order structure size with the additive pinning approach. On the upper panels we depict the snapshots of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units, while the lower panels show the hyperedge-based local order parameters. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. Panels a) and b) report the simulations performed on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes, panels c) and d) on a $4$-hyperring (i.e., $5$-body interactions) of $272$ nodes, panels e) and f) on a $5$-hyperring (i.e., $6$-body interactions) of $340$ nodes and panels g) and h) on a $6$-hyperring (i.e., $7$-body interactions) of $408$ nodes. The number of nodes is chosen so that each hyperring is made of $68$ hyperedges. Pinning control is applied to $N_{p}=18$ nodes as in Fig. 7, i.e., every two junction nodes, to all the structures and for $t_{p}=100$ time units. This means that the pinned nodes are $\simeq 8.8\%$ of the total nodes in the $3$-hyperring, $\simeq 6.6\%$ of the total nodes in the $4$-hyperring, $\simeq 5.3\%$ of the total nodes in the $5$-hyperring and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring. The model parameters are $\alpha=1$ and $\omega=1$, the coupling strength is $\varepsilon=0.01$, except for the $6$-hyperring where $\varepsilon=0.1$, and the parameters $\lambda_{i_{p}}$ are drawn from a uniform distribution in the interval $[0,2]$ and are the same for all the simulations.
@@
In Fig. 8, we show the results obtained with such control scheme on $d$-hyperrings with $d=3,4,5,6$, i.e., $4$-, $5$-, $6$- and $7$-body interactions, where we have fixed the number of hyperedges. Indeed, we can observe that we obtain a chimera state by inducing a large region of incoherence (more than half of the nodes) with a control that involves only a small fraction of the nodes. Moreover, the pinned nodes $\mathcal{I}_{p}=\{1,\ldots,N_{p}\}$ are kept constant for every structure, meaning that $N_{p}$ does not scale with the number of nodes, but rather with the number of hyperedges, which allows to control large structures with only a handful of nodes. The pinned nodes are $\simeq 8.8\%$ of the total nodes for the $3$-hyperring (panels a) and b)), $\simeq 6.6\%$ of the total nodes for the $4$-hyperring (panels c) and d)), $\simeq 5.3\%$ of the total nodes in the $5$-hyperring (panels e) and f)) and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring (panels g) and h)). In the latter case, our pinning scheme does not work as well as for the other structures and the coupling strength needs to be significantly increased in order to observe a chimera state. However, we can observe from Fig. 8h) that the chimera is not anymore stable and the front of incoherence enlarges. Let us stress that stable chimera states can be easily obtained also in $6$-hyperrings by reducing the distance between the controlled nodes, as in the previous section, instead of the pinning scheme of Fig. 7.
@@
Let us note that, for the pinning scheme of Fig. 7, the parameters $\lambda_{i_{p}}$ need to be all positive or all negative in order for this pinning scheme to yield persistent chimera states. When such random inputs are drawn in a symmetric interval with respect to the 00, e.g., $[-2,2]$ as in Fig. 3, a chimera states forms but turns into a weak chimera-like state at about $400$ time units. On the other hand, when the control inputs have the same sign, e.g., $[0,2]$ as in Fig. 8, the chimera is persistent until about $4000$ time units, i.e., $10$ times longer, before turning into a weak chimera-like state.
@@
In Appendix B, we show the results for the parametric pinning, that are analogous except for the case of $6$-hyperring, where only weak chimera-like states emerge when pinning every other junction node, even with stronger couplings. The fact that our results are robust with respect to different control approaches is a good indication of the pervasiveness of the phenomenon.
@@ IV Heuristic interpretation through phase reduction theory
In this section we will give a heuristic interpretation of the results based on the phase reduction approach nakao2016phase ; kuramoto2019concept , which consists of reducing a given oscillatory system to a phase, i.e., Kuramoto-type, model kuramoto1975 . In a nutshell, starting from a system of $N$ highly dimensional units in a limit cycle regime, e.g.,
@@
under given assumptions (see nakao2016phase for details), we can reduce to a system of $N$ interacting phase oscillators of the form
$\dot{\vartheta}_{i}=\omega_{i}+\varepsilon\sum_{j=1}^{N}A_{ij}\vec{Z}(\vartheta_{i})\cdot\vec{G}_{ij}(\vartheta_{j},\vartheta_{i}),$ (20)
where $\omega_{i}$ is the frequency of the $i$-th oscillator and $\vec{Z}$ is the phase sensitivity function. The key in the reduction process is to find an expression for $\vec{Z}$, which is analytical only in some specific cases, among which the Stuart-Landau model nakao2016phase and weakly nonlinear oscillators leon2023analytical , while it needs to be obtained numerically for general oscillators.
The phase reduction theory has been applied also to systems with higher-order interactions leon2025theory ; ashwin2016hopf ; leon2019phase ; leon2024higher , obtaining higher-order versions of the Kuramoto model, which exhibit much richer behaviors than the pairwise one. Indeed, the first evidence of higher-order-induced exotic behaviors, which triggered the interest of the community towards this new framework, has come from higher-order phase models (although not derived through phase reduction) tanaka2011multistable ; bick2016chaos ; skardal2019abrupt ; skardal2019higher ; millan2020explosive . In what follows, we will apply the phase reduction to our model on a $3$-hyperring and on the corresponding clique-projected network, Eqs. (6) and (7) respectively, to give an intuition of why it is easier to induce chimera behavior via pinning when the topology is higher-order.
For the Stuart-Landau model, the phase sensitivity function is $\vec{Z}(\vartheta)=(-\sin(\vartheta),\cos(\vartheta))^{\top}$ nakao2016phase . Let us consider system (6) in polar coordinates, i.e., $\vec{X}_{i}=(\cos(\vartheta_{i}),\sin(\vartheta_{i}))$, and proceed with the reduction by computing the following
@@
$-\alpha\cos(\vartheta_{i})\sin(\vartheta_{i})+\omega\sin^{2}(\vartheta_{i})+\cos(\vartheta_{i})\sin(\vartheta_{i})+\omega\cos^{2}(\vartheta_{i})+\alpha\cos(\vartheta_{i})\sin(\vartheta_{i})$
$-\cos(\vartheta_{i})\sin(\vartheta_{i})+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}\cos^{3}(\vartheta_{i})\sin(\vartheta_{i})-$
$\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\sin(\vartheta_{i})+\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\cos(\vartheta_{i})-\cos^{4}(\vartheta_{i})\Big{)},$
@@
$\dot{\vartheta}_{i}=\omega+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Phi(\vartheta_{i},\vartheta_{j_{1}},\vartheta_{j_{2}},\vartheta_{j_{3}}),$ (21)
@@
$\Phi(\vartheta_{i},\vartheta_{j_{1}},\vartheta_{j_{2}},\vartheta_{j_{3}})$ $=$ $\cos^{3}(\vartheta_{i})\sin(\vartheta_{i})-\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\sin(\vartheta_{i})+$ (22)
$\cos(\vartheta_{j_{1}})\cos(\vartheta_{j_{2}})\cos(\vartheta_{j_{3}})\cos(\vartheta_{i})-\cos^{4}(\vartheta_{i}).$
@@
$\dot{\vartheta}_{i}=\omega+\varepsilon\sum_{j}^{N}A_{ij}\Psi(\vartheta_{i},\vartheta_{j}),$ (23)
@@
Through the averaging method nakao2016phase , the $4$-body coupling (22) can be approximated as
@@
The coupling given by Eq. (26) steers the system towards synchronization, because $\Psi(\vartheta_{i},\vartheta_{j})$ has form of the well-known Kuramoto-Sakaguchi coupling, i.e., $\sin(\vartheta_{j}-\vartheta_{i}+\alpha)$, which is known to be attractive for $|\alpha|<\frac{\pi}{2}$ Kuramoto . On the other hand, Eq. (25) allows for a much richer dynamics, given that there are many more combinations of the phases and the coefficients such that the coupling term vanishes, as it is the case of the higher-order Kuramoto model leon2024higher . This fact gives an intuition not only of the much richer dynamics observed when higher-order interactions are present tanaka2011multistable ; skardal2019higher ; millan2020explosive ; leon2024higher ; skardal2019abrupt , but also of why higher-order systems favor the presence of chimera states and it is much easier to induce such behavior via pinning, compared to the pairwise case. Moreover, given the form of the higher-order coupling terms, once such state is achieved, it is more difficult for the higher-order system to steer towards synchronization, which provides a qualitative explanation of why the chimera states are also persistent. Let us remark that the first intuition of this behavior was given in zhang2024deeper , where it was shown, for the higher-order Kuramoto model, that, when the system leaves the attraction basin of the synchronous state, it is more difficult to synchronize again because higher-order interactions cause a shrinking of such attraction basin, which becomes smaller. A similar conclusion has been provided also in von2024higher . In our case, the control pushes the system away from the synchronous solution creating a chimera state and the higher-order interactions favor the persistence of this state.
@@
Figure 9: Pinning induces phase chimera states for the reduced phase model (21) on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$. The frequency is $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of Figs. 3 and 4.
@@
Figure 10: Pinning for the reduced phase model (23) on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 9. Panel a) depicts the whole time series of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the cosine of the phases $\cos(\vartheta_{i}(t))$ with $i=1,...,N$. The frequency is $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of Figs. 3 and 4.
@@
As a further corroboration of the results shown in this work and of the correctness of our phase reduction approach, let us conclude this section by applying our pinning protocols to the reduced phase models and show that the outcome confirms our claims. Before proceeding with the implementation of the pinning protocol, let us observe that for a phase model, such as that of Eq. (21), the additive and parametric protocols are equivalent. In fact, the uncoupled oscillators are described by the sole frequency $\omega$, hence, an additional input (i.e., additive pinning) or a modification of the latter (parametric pinning) are the same operation. It is important to note that this does not mean that the two pinning protocols are always equivalent, but, when the coupling is small enough for the phase reduction to be valid, the approximated phase model obtained after averaging reacts in the same way for both protocols. In Figs. 9 and 10, we show the result of the pinning procedure on a $3$-hyperring and its corresponding clique-projected network, respectively. The results are analogous to those of the non-reduced model: while the control induces a chimera state in the higher-order phase reduced model, no chimeras emerge in the pairwise setting. This was expected, given that we are dealing with phase chimeras and, hence, the phase model should fully reproduce such patterns. Again, note that the phase reduction is an approximation, and the phase reduced model does not behave exactly as the full model, but there may be small differences [^5] These differences can be particularly relevant close to bifurcation points leon2019phase ; bick2024higher , but this is not the case in the setting we consider here..